OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 2x + 3y = 12 and 3x + 2y = 13 (3, 2) (-3, 2) (-3, -2) no solution infinitely many solutions

OpenStudy (anonymous):

in this case it is easiest just to check the solutions provided. plug in the numbers and see if they work

OpenStudy (anonymous):

only real chance is for the first one. try \[2\times 3+3\times 2=6+6=12\] \[3\times 3+2\times 2=9+4=13\] so first one it is

OpenStudy (radha):

multiply the first eqn with 3...and the second one with 2...then eliminate...!

OpenStudy (radha):

you will get 6x+9y=36.......6x+4y=26 now by solving by subtracting 2nd eqn from 1st....we get y=2 and x=3....so first option is correct